OpenStudy (anonymous):

Solve the system of equations. x + 3y = -1 2x + 2y = 6 (-4, 1) (2, -1) (4, -1) (5, -2)

OpenStudy (anonymous):

Given the following system of eqn, you could either use the substitution or elimination method \(\begin{align}x + 3y &= -1 \\2x+2y &= 6\end{align}\) Let's go in favour of the first We get an expression with respect to \(x\) \(x = -3y - 1\) Substiture value of \(x\) into the second given eqn and solve for \(y\)

OpenStudy (anonymous):

add 3 to each side , right?

OpenStudy (anonymous):

Nope, I meant where you see \(x\) put \(-3y-1\) right there

OpenStudy (anonymous):

so -3y-1

OpenStudy (anonymous):

Like this\(\ldots ~2(-3y-1) +2y = 6\)

OpenStudy (anonymous):

Expand that and solve for \(y\). Can you attempt that?

OpenStudy (anonymous):

now i do the 2 sides?

OpenStudy (anonymous):

Show me what you mean.

OpenStudy (anonymous):

+3y +1

OpenStudy (anonymous):

i got x=-3/4

OpenStudy (anonymous):

That's incorrect \(\bigotimes\)

OpenStudy (anonymous):

D

OpenStudy (anonymous):

5,-2

OpenStudy (anonymous):

You're only guessing. :(

OpenStudy (anonymous):

i just got y=-2 from doing the equation

OpenStudy (anonymous):

Expanding this: \(2(-3y-1) +2y = 6\) gives \(-6y-2 +2y=6 \longrightarrow -4y = 8 \longrightarrow y=-2\) Substitute \(y= -2\) in \(x + 3y = -1\), you'll get \(x + 3(-2) = -1 \longrightarrow x - 6 = -1 \longrightarrow x = 5\) So, you get your guessed answer, \(\left(5, -2\right)\)

OpenStudy (anonymous):

It wasnt guessed.

OpenStudy (anonymous):

"i just got y=-2 from doing the equation" ??? \(\longleftarrow\) I don't think so, simply because you couldn't show your work. But it's okay